OpenStudy (anonymous):

how do you find the volume of a rectangular prisim whose lenght is 2x+3 and width is 4w and height is x. i know the formula V=Area of base*height. How do i solve it with these variables. it does not give the measure of the volume in the problem. it just says use algebra to express the volume of the right rectangular prisim. so how will i do that step by step.?

OpenStudy (mertsj):

V = Area of base times height Area of base = (2x+3)(4w) Height = x V=x(4w)(2x+3)=8wx^2+12xw Although I suspect you meant to write that the width is 4x

OpenStudy (anonymous):

so the final answer would be V=8wx^2+12xw so i really can't have an offical number for it is that correct?

OpenStudy (mertsj):

That is correct. Are you sure that the width is 4w or is it 4x?

OpenStudy (anonymous):

its 4w

OpenStudy (mertsj):

ok. That's your answer then.

OpenStudy (anonymous):

okay thanks!